OpenStudy (anonymous):

The graph of which equation has a vertex of (2, 5)? (3 points) Answer y = 3(x - 2)2 + 5 y = 5(x - 2)2 + 1 y = 3(x + 2)2 + 5 y = 2(x + 5)2 + 2

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

heyhey

hartnn (hartnn):

general equation : y= a(x-h)^2 +k where vertex is (h,k) and bdw, hi :)

OpenStudy (anonymous):

:) my answer is C

hartnn (hartnn):

nopes.

hartnn (hartnn):

(x-h) for C, (x+2) ---> h=-2

OpenStudy (anonymous):

So it's either A or B

hartnn (hartnn):

in B. k=1

OpenStudy (anonymous):

:(((((((((((((

hartnn (hartnn):

? so, what remains ?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

;)

hartnn (hartnn):

yes.

OpenStudy (anonymous):

thank you!!!!!

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]

OpenStudy (anonymous):

wow thats a cool "welcome"!!!!!!!!!!!!!!

hartnn (hartnn):

;)